OpenStudy (anonymous):

Factor completely: 36x2 - 49

OpenStudy (anonymous):

oops the 2 is an exponent

OpenStudy (saifoo.khan):

"differences of squares"

OpenStudy (anonymous):

36x^2=(6x)^2 and 49=7^2 then just apply the difference of square

OpenStudy (anonymous):

(6x-7)(6x+7)

OpenStudy (theviper):

\[\Huge{\color{red}{36x^2-49}}\]\[\LARGE{\color{green}{\implies(6x)^2=(7)^2}}\]\[\LARGE{\color{green}{\implies(6x)^2-(7)^2}}\]\[\LARGE{\color{green}{\implies(6x+7)(6x-7)}}\] Remember,\[\Huge{{\star \star \star}\color{orange}{(a^2-b^2)=(a+b)(a-b)}}\]

OpenStudy (saifoo.khan):

lol. wow.

OpenStudy (theviper):

Thanx!